# TRIM: Token Reduction and Inference Modeling for Cost-Effective Language Generation

### v1

@@ Abstract
The inference cost of Large Language Models (LLMs) is a significant challenge due to their computational demands, specially on tasks requiring long outputs. However, natural language often contains redundancy, which presents an opportunity for optimization. We have observed that LLMs can generate distilled language—concise outputs that retain essential meaning, when prompted appropriately. We propose a framework for saving computational cost, in which a shorter distilled output from the LLM is reconstructed into a full narrative by a smaller model with lower inference costs. Our experiments show promising results, particularly in general knowledge domains with 20.58% saved tokens on average with tiny decrease in evaluation metrics, hinting that this approach can effectively balance efficiency and accuracy in language processing tasks.
@@ 1 Introduction
In this work, the main contribution is the proposal of a novel technique for reducing the inference costs of LLMs through the suppression and later-on reconstruction of function words. This approach is orthogonal to other optimization techniques, and could be applicable as LLMs continue to grow in size and capabilities. We also propose an algorithm to check and define the applicability of this technique in different domains, selecting the most proper function words set, and analyzing the lose in performance as the percentage of saved tokens increases. Additionally, we provide an experimental evaluation in the context of general knowledge question-answering. In this setting, the output length is more relevant than the input query, making the applicability of our technique more evident. The results show that even suppressing around 18% of the full text, models that are hundreds of times smaller can reconstruct distilled text without losing significant text quality.
@@ 3 Exploratory Research
4 Semantic Relevance Distribution Of Words In The English Language
@@ 5 Cost-Saving Pipeline
The exploratory research showed the feasibility of the initial idea. Now, we cover the development and specification of the approach. We propose a pipeline for wrapping the use of generative LLMs (both closed and open) to save costs during output generation through the philosophy of omitting the generation of inferable function words, and reconstructing the full text using a much smaller and optimized language model later on. This pipeline can be seen in Figure 1. It can be defined by the following elements:
@@
Figure 1: Diagram of the proposed methodology against a vanilla technique.
@@ 6 Datasets & Models
Second, we used a synthetic knowledge question-answering dataset we have created for this project. We required an evaluation that aligns with the proposed methodology where the reconstruction model reconstructs distilled text generated by an LLM. For this purpose, we generated 9,000 knowledge questions on different topics. For the generation, we took 30 general knowledge topics, 20 concepts on each of them and 5 keywords associated to each concept, creating various questions from templates regarding explanation, summarization and concept comparison. See supplementary material (Section A.2) for specifications, and an example of these questions.
@@ 7 Evaluation
For the evaluation, we thought it was necessary to find a specific metric that would evaluate the recovery of the terms from the inferable set. The common NLP metrics focus on comparing two pieces of text that can come in different sizes (i.e., word, sentence, paragraph…), but they do not offer the versatility of focusing only on a specific set of terms according to their exact position, covering the cases of omission, substitution, matching or insertion. The Needleman-Wunsch algorithm, used in bio-informatics to align protein or nucleotide sequences, offers these desired features Needleman and Wunsch (1970), so we adapted it to align the paragraphs of the original and distilled text to effectively maintain the alignment over scenarios where there are missing or extra terms. After applying the algorithm, we can compare each of the positions in the alignment and cover the cases mentioned above, obtaining precision, recall and F1 values with regard to the inferable set. From now on, these metrics will be referred as $\Theta$ metrics.
@@ 9 Conclusion
In this work we proposed an algorithm for the identification of a set of easily-inferable terms based on the context according to the desired cost-savings by using masking together with language models. We have proposed a pipeline that allows cost savings in LLM-generation by omitting the generation of terms that do not contribute to the semantic meaning, and can be easily inferred a posteriori with a smaller language model. We have evaluated our approach in the question-answering knowledge domain, where we obtained a subset of functions words that are easily inferable. We have trained 4 different language models in the task of reconstructing a distilled answer. These models obtain very good performance across all metrics without noticeable losses, while saving 20.58% tokens with respect to the original full answer. We also showed that, while the output generation is more expensive than processing input prompts, in the vast majority of cases it will be better of to process knowledge answers through the cost-saving pipeline.
@@ 10 Limitations
Finally, another relevant limitation is that be have used automated metrics such as SacreBLEU, ROUGE, and perplexity. This provide valuable insights into grammatical similarity, semantic coherence, and fluency, but they may not fully capture the nuances of human comprehension and acceptability Warstadt (2019), especially in reconstructed texts. A future human evaluation would be desired to assess subtle aspects like readability, naturalness, and the overall quality of the reconstructed responses, which metrics alone may overlook. Including human evaluators would enable a more comprehensive understanding of the method’s effectiveness, particularly in scenarios where slight changes in phrasing or word choice could impact clarity and user satisfaction.

### v2

@@ Abstract
The inference cost of Large Language Models (LLMs) is a significant challenge due to their computational demands, specially on tasks requiring long outputs. However, natural language often contains redundancy, which presents an opportunity for optimization. We have observed that LLMs can generate distilled language—concise outputs that retain essential meaning, when prompted appropriately. We propose TRIM, a pipeline for saving computational cost in which a shorter distilled output from the LLM is reconstructed into a full narrative by a smaller model with lower inference costs. Our experiments show promising results, particularly in general knowledge domains with 20.58% saved tokens on average with tiny decrease in evaluation metrics, hinting that this approach can effectively balance efficiency and accuracy in language processing tasks.
@@ 1 Introduction
In this work, the main contribution is the proposal of a novel pipeline for reducing the inference costs of LLMs through the suppression and later-on reconstruction of function words. This approach is orthogonal to other optimization techniques, and could be applicable as LLMs continue to grow in size and capabilities. We also propose an algorithm to check and define the applicability of this technique in different domains, selecting the most proper function words set, and analyzing the lose in performance as the percentage of saved tokens increases. Additionally, we provide an experimental evaluation in the context of general knowledge question-answering. In this setting, the output length is more relevant than the input query, making the applicability of our technique more evident. The results show that even suppressing around one fifth of the full text, models that are hundreds of times smaller can reconstruct distilled text without losing significant text quality.
@@ 3 Exploratory Research
4 Semantic Relevance Distribution Of Words
@@ 5 Cost-Saving Pipeline
The exploratory research showed the feasibility of the initial idea. Now, we cover the development and specification of the approach. We propose TRIM (Token Reduction and Inference Modeling pipeline for cost-effective language generation), a pipeline for wrapping the use of generative LLMs (both closed and open) to save costs during output generation through the philosophy of omitting the generation of inferable function words, and reconstructing the full text using a much smaller and optimized language model later on. This pipeline can be seen in Figure 1. It can be defined by the following elements:
@@
Figure 1: Diagram of the proposed cost-saving pipeline with examples
@@ 6 Datasets & Models
Second, we used a synthetic knowledge question-answering dataset we have created for this project. We required an evaluation that aligns with the proposed TRIM pipeline where the reconstruction model reconstructs distilled text generated by an LLM. For this purpose, we generated 9,000 knowledge questions on different topics. For the generation, we took 30 general knowledge topics, 20 concepts on each of them and 5 keywords associated to each concept, creating various questions from templates regarding explanation, summarization and concept comparison. See supplementary material (Section A.2) for specifications, and an example of these questions.
@@ 7 Evaluation
For the evaluation, we thought it was necessary to find a specific metric that would evaluate the recovery of the terms from the inferable set. The common NLP metrics focus on comparing two pieces of text that can come in different sizes (i.e., word, sentence, paragraph…), but they do not offer the versatility of focusing only on a specific set of terms according to their exact position, covering the cases of omission, substitution, matching or insertion. The Needleman-Wunsch algorithm, used in bio-informatics to align protein or nucleotide sequences, offers these desired features Needleman and Wunsch (1970), so we adapted it to align the paragraphs of the original and reconstructed text to effectively maintain the alignment over scenarios where there are missing or extra terms. After applying the algorithm, we can compare each of the positions in the alignment and cover the cases mentioned above, obtaining precision, recall and F1 values with regard to the inferable set. From now on, these metrics will be referred as $\Theta$ metrics.
@@ 9 Conclusion
In this work we proposed an algorithm for the identification of a set of easily-inferable terms based on the context according to the desired cost-savings by using masking together with language models. We have proposed TRIM, a pipeline that allows cost savings in LLM-generation by omitting the generation of terms that do not contribute to the semantic meaning, and can be easily inferred a posteriori with a smaller language model. We have evaluated our approach in the question-answering knowledge domain, where we obtained a subset of functions words that are easily inferable. We have trained 4 different language models in the task of reconstructing a distilled answer. These models obtain very good performance across all metrics without noticeable losses, while saving 20.58% tokens with respect to the original full answer. We also showed that, while the output generation is more expensive than processing input prompts, in the vast majority of cases it will be better of to process knowledge answers through the cost-saving pipeline.
@@ 10 Limitations
Finally, another relevant limitation is that we have used automated metrics such as SacreBLEU, ROUGE, and perplexity. This provide valuable insights into grammatical similarity, semantic coherence, and fluency, but they may not fully capture the nuances of human comprehension and acceptability Warstadt (2019), especially in reconstructed texts. A future human evaluation would be desired to assess subtle aspects like readability, naturalness, and the overall quality of the reconstructed responses, which metrics alone may overlook. Including human evaluators would enable a more comprehensive understanding of the method’s effectiveness, particularly in scenarios where slight changes in phrasing or word choice could impact clarity and user satisfaction.
@@ A.3 Consistency among responses in LLMs
A.4 Aligning Function Words and $\Theta$-metrics
$\Theta$-metrics are the metrics designed and used in our evaluation with the purpose of considering only the performance over the function words prediction or reconstruction. Consider the following example:
  • Original: The history of art is the fascinating subject of human culture.
  • Reconstructed: History of the art is a fascinating subject of human culture.
After removing all non-function words and aligning the resulting sequences with the Needleman-Wunsch algorithm Needleman and Wunsch (1970), we can compute matches and errors as in Table 7. In terms of binary classification result, the matches correspond to the true positives. The false positives are the insertions plus the substitutions. The false negative are the omission plus the substitutions. Then, precision, recall and F1 score are computed in the standard way.
Original The of the of
Reconstructed of the a of
Result omission match insertion substitution match
Table 7: Example of aligned function words and matching result.